OpenStudy (anonymous):

the area of a rectangular garden is 234 square meters. the width is 4 meters longer than one-half of the length. find the length and the width of the garden.

OpenStudy (whpalmer4):

\[A=l*w\] We know that \(w = 4+\frac{1}{2}l\) so we can substitute that into the equation: \[A = l*w = l*(4+\frac{1}{2}l) = 234\]Solve for \(l\), then plug \(l\) into the equation for \(w\)

OpenStudy (whpalmer4):

Note that you'll get two solutions for \(l\), but only one of them makes sense in the context of this problem.

OpenStudy (anonymous):

 this might help you